OpenStudy (anonymous):

Solve X^2 + x -30 X -5

OpenStudy (anonymous):

x=-5? -5^2-5-30 25-5-20= 20-30=-10

OpenStudy (anonymous):

Thanks!!(:

OpenStudy (compassionate):

Hi, This is a trinomial because it has three terms. A trinomial can be simplified into a binomial. A binomial akes the form (x -+ y)(x -+ z) You want to start off by solving what you know! You know that the first term MUST be x to the second power. So, that's a good place to start! \[(x \pm ?)(x \pm ?) \] You are given the expression: \[ x^2 + x -30 \] When added together, what will produce x and when multiplied produce -30? Did you say 6 and -5? Correct! So we have: \[(x + 6)(x - 5)\]Now, al that is being divded by x-5. So lets plug those numbers in: \[\frac{ (x + 6)(x - 5) }{ x-5 }\]x-5 and x-5 cancel out! Leaving you with x + 6! OpenStudy Ambassador: Compassionate

OpenStudy (anonymous):

thanks!(: